# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

### v1

@@ Introduction
We evaluate the performance of our method using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRation, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
Main Results
@@
We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques. The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace. The second baseline, termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
@@
The results of LLaMA-3.1-8B-Instruct are summarized in Table 1. As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens. Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning. In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.55 at Ratio = 0.5) while maintaining robust reasoning performance. In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2). In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively. These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
We also provide the results of Qwen2.5‐Instruct series models as Table 2. Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression. For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31). As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%). Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%. Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget.
@@
Figure 4: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
@@
Table 2: Experimental results on the Qwen2.5-Instruct series. We also provide Accuracy, Tokens and ActRatio.
Error Propagation via Controlled Injection
To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conduct a systematic error-injection study to test the “cascading failure” hypothesis. Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark. For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of correct CoT sequences. Then, we evaluate how these perturbations affect the overall reasoning process.
The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation. By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position of such errors affects the final reasoning outcome.
@@
Table 3 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B. When injecting a numerical error, the accuracy drops are relatively uniform across positions: such as a miscalculation at step 4 a drop of 61.70%. In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points. Replacement of an operator in the final step (4/4) inflicts a reduction of 80.80%. These results confirm that CoT reasoning exhibits “late-stage fragility” particularly to logical perturbations. More details can be found in the appendix.
Analysis
Length Budget Analysis
To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure  4, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (0.5–1.0). While in the second, we retain the original LLM budget (e.g., 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8, and 0.9, we observe absolute gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT within the same budgetary constraints.
The Late-Stage Fragility of CoT
Our findings reveal a phenomenon that challenges the traditional cascading failure model. To illustrate this conceptually, Figure  5 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps. However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
@@
Figure 5: A illustration of “Late-Stage Fragility”. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Case Study: ASCoT in Action
To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure  3. The problem asks for the total number of apples Bob picked over nine days.
@@
All/Err Type Ori Acc Fin Acc Drop Ratio
2/2 numeric 98.85 38.93 59.92
3/3 96.19 35.72 60.47
4/4 95.75 34.05 61.70
2/2 symbolic 98.85 29.18 69.67
3/3 96.19 20.42 75.77
4/4 95.75 14.95 80.80
Table 3: Error-injection results on GSM8K, using Qwen2.5-Instruct 7B. “All/Err” denotes the total number of reasoning steps and the step which a fault was injected. “Type” indicates the category of the injected error (numeric or symbolic). “Ori Acc” is the model’s accuracy before error-injection. “Fin Acc” is the accuracy after injection. “Drop Ratio” quantifies the relative decrease in accuracy.
Related Work
LLM Reasoning & Efficient Thinking
@@
The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs (Marconato et al. 2023; Ding et al. 2024; Liu et al. 2024d). CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks (Team 2024).
@@
These efficiency-focused methods can be broadly categorized: Prompt-based strategies (Xu et al. 2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget (Han et al. 2024). Inference-time dynamic reasoning (Ding et al. 2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation. CoT Distillation (Wadhwa, Amir, and Wallace 2024; Wang et al. 2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
CoT Compression and Pruning
@@
The verbosity of CoT (Wingate, Shoeybi, and Sorensen 2022), while beneficial for reasoning, leads to high inference latency and computational costs. This has spurred research into CoT compression and pruning (Chuang et al. 2024; Xuefei et al. 2023), which aims to eliminate redundant information without degrading reasoning accuracy. This issue is often termed the “overthinking phenomenon,” (Kumar et al. 2025; Chen et al. 2024b) where models generate unnecessarily long chains that can even introduce errors.
@@
To address this bottleneck, researchers have explored various techniques. Token-level pruning (Fu et al. 2024; Ma, Fang, and Wang 2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts. Other approaches operate at a more structural level. Prune-on-Logic (Liu et al. 2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score. Simpler methods, such as direct truncation of the reasoning chain (Wang et al. 2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
Self-correction Mechanism
@@
The capacity for self-correction is a critical frontier (Pan et al. 2023) for enhancing the reliability of LLMs. The general paradigm (Kamoi et al. 2024; Yang et al. 2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
@@
Approaches diverge based on the source of the feedback. Intrinsic self-correction (Li et al. 2024; Liu et al. 2024a) relies on the model’s own capabilities to critique and improve its output without external tools (Cao, Wang, and Xiong 2025; Tan et al. 2025). However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data (Liu et al. 2024b; Anantaprayoon, Kaneko, and Okazaki 2025). Extrinsic self-correction incorporates feedback from external sources. This involves using a verifier model to assess the correctness of each step (Kim et al. 2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
Conclusion

### v3

@@ 4.1. Experimental Setup
Evaluation Metrics We evaluate the performance of ASCoT using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
4.2. Main Results
@@
We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques. The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace. The second termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
@@
The results of LLaMA-3.1-8B-Instruct are summarized in Table 1. As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens. Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning. In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.79 at Ratio = 0.8 on GSM8K and ActRatio = 0.61 at Ratio = 0.6 on MATH-500) while maintaining robust reasoning performance. In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2). In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively. These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
Figure 4. Performance of ASCoT on the Qwen2.5-Instruct Series. The results for the 3B, 7B, and 14B models are shown under various compression ratios and compared against the original, uncompressed baseline.
@@
Figure 5. Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
We also provide the results of Qwen2.5‐Instruct series models as Table 2. Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression. For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31). As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%). Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%. Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget. The detailed experimental results can be found in the appendix.
Table 2. Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We also provide Accuracy, Tokens and ActRatio.
@@
Figure 4 demonstrates that ASCoT consistently achieves substantial CoT compression across all tested Qwen2.5-Instruct series models (3B, 7B, and 14B). For the 3B model, a moderate compression that reduces token usage by nearly 17% (from 314.87 to 262.31 tokens) incurs only a negligible 0.5-point drop in accuracy. This trend is further amplified in the larger models. With the 7B variant, a 15% token reduction results in a minimal 0.1-point accuracy drop, and a nearly 50% reduction in tokens still maintains accuracy above 86%. Notably, the 14B model showcases ASCoT’s exceptional ability to enhance efficiency. Even when halving the original token budget (a 50% reduction to 158.21 tokens), the accuracy drop is a mere 1.5 percentage points. These results validate ASCoT’s effectiveness in striking a favorable balance between computational efficiency and reasoning fidelity, demonstrating its robustness and scalability across different model sizes.
4.3. Error Propagation via Controlled Injection
To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conducted a comprehensive and systematic error-injection study to rigorously test the “cascading failure” hypothesis. Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark. For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of the correct CoT sequences. Then, we assess and evaluate how these perturbations affect the overall reasoning process.
The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation. By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position and type of such errors affects the final reasoning outcome.
@@ 4.3. Error Propagation via Controlled Injection
Table 4 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B. For numeric errors, the accuracy drops gradually across positions, with a miscalculation at step 2 causing a decrease of 14.64%, at step 3 a 29.56% drop, and at step 4 a significant reduction of 51.69%. In contrast, symbolic errors induce more severe performance degradation, especially when injected later in the reasoning process. A symbolic error at step 2 results in a large drop of 43.97%,. However, injecting a symbolic error at step 3 (3/3) results in an accuracy reduction of 56.78%. These results emphasize the susceptibility of CoT reasoning to errors, particularly with symbolic mistakes, which cause a more pronounced decline in performance as the reasoning progresses. Table 3 reports initial accuracy before error injection. All the above data are from dataset GSM8K.
Table 3. Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
Steps Correct Total Accuracy
0 615 660 93.18
1 1 1 100.00
2 777 786 98.85
3 2044 2125 96.19
4 1936 2022 95.75
5 1110 1180 94.07
6 440 484 90.91
7 117 137 85.40
8 39 58 67.24
9 4 7 57.14
10 11 11 100.00
11 2 2 100.00
4.4. Analysis
Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure 5, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0). In the second, we retain the original LLM budget (e.g., 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budgetary constraints.
The Late-Stage Fragility of CoT Our findings reveal a phenomenon that challenges the traditional cascading failure model. To illustrate this conceptually, Figure 6 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process. An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps. However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
@@ 4.4. Analysis
Figure 6. A illustration of “Late-Stage Fragility”. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure 3. The problem asks for the total number of apples Bob picked over nine days.
@@
Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
5. Related Work
5.1. LLM Reasoning & Efficient Thinking
@@
The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs Marconato et al. (2023); Ding et al. (2024); Liu et al. (2024a). CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks Team (2024).
@@
These efficiency-focused methods can be broadly categorized: Prompt-based strategies Xu et al. (2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget Han et al. (2024). Inference-time dynamic reasoning Ding et al. (2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation. CoT Distillation Wadhwa et al. (2024); Wang et al. (2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
Table 4. Error-injection results on GSM8K, using Qwen2.5-Instruct 7B. “All/Err” denotes the total number of reasoning steps and the step which a fault was injected. “Type” indicates the category of the injected error. “Ori Acc” is the model’s accuracy before error-injection. “Fin Acc” is the accuracy after injection. “Accu Drop(abs. points)” quantifies the relative decrease in accuracy.
All/Err Type Ori Acc Fin Acc Accu Drop
4/2 numeric 95.75 81.11 14.64
4/3 95.75 66.19 29.56
4/4 95.75 44.06 51.69
2/2 symbolic 98.85 54.88 43.97
3/2 96.19 78.28 17.91
3/3 96.19 39.41 56.78
5.2. CoT Compression and Pruning
@@
The verbosity of CoT Wingate et al. (2022), while beneficial for reasoning, leads to high inference latency and computational costs. This has spurred research into CoT compression and pruning Chuang et al. (2024); Xuefei et al. (2023), which aims to eliminate redundant information without degrading reasoning accuracy. This issue is often termed the “overthinking phenomenon,” Kumar et al. (2025); Chen et al. (2024a) where models generate unnecessarily long chains that can even introduce errors.
@@
To address this bottleneck, researchers have explored various techniques. Token-level pruning Fu et al. (2024); Ma et al. (2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts. Other approaches operate at a more structural level. Prune-on-Logic Liu et al. (2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score. Simpler methods, such as direct truncation of the reasoning chain Wang et al. (2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
5.3. Self-correction Mechanism
@@
The capacity for self-correction is a critical frontier Pan et al. (2023) for enhancing the reliability of LLMs. The general paradigm Kamoi et al. (2024); Yang et al. (2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
@@
Approaches diverge based on the source of the feedback. Intrinsic self-correction Li et al. (2024); Liu et al. (2024c) relies on the model’s own capabilities to critique and improve its output without external tools Cao et al. (2025); Tan et al. (2025). However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data Liu et al. (2024b); Anantaprayoon et al. (2025). Extrinsic self-correction incorporates feedback from external sources. This involves using a verifier model to assess the correctness of each step Kim et al. (2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
6. Limitations
While ASCoT demonstrates promising performance, we acknowledge that there still exist several limitations that define avenues for future research. The efficacy of the framework is inherently dependent on the self-evaluation capabilities of its base Large Language Model (LLM); if a model cannot fundamentally recognize its own fallacies, the utility of the AVM and MSCE modules is consequently constrained. Furthermore, our current evaluation is focused on numeric and symbolic errors within the specific domain of mathematical reasoning. Therefore, the generalization of ASCoT to other error modalities—such as factual, commonsense, or complex logical errors—and its transferability to diverse long-chain reasoning tasks like code generation, legal analysis, or scientific summarization, remain important open questions requiring further study.
7. Conclusion
@@
Appendix A Appendix
This document provides supplementary material, including pseudocode for the ASCoT mechanism, experimental setup, error-injection experiments, and results analysis. It also covers model configurations, hyperparameters used in the study. Additionally, the document discusses the model’s late-stage fragility and outlines current limitations and future directions.
A.1. Experimental Setup
To clearly elucidate the algorithmic pipeline of the ASCoT methodology, we provide its pseudocode. The algorithm commences upon receiving a problem query, which is first processed by the target Large Language Model (LLM) to generate an initial CoT solution, denoted as $CoT_{initial}$. This initial output is subsequently passed to the IRM module for token pruning based on a predefined compression ratio $\gamma$, thereby enhancing efficiency. The compressed CoT then undergoes a step-wise risk assessment by the AVM module. For each step $t_{k}$, the AVM computes a Risk Score $R(t_{k})$, which integrates a multi-dimensional quality assessment $Q(t_{k})$ with a Positional Impact score $I(k)$. If this score exceeds a predefined threshold $\tau$, the step is flagged as high-risk and is routed to the MSCE module for remediation. The MSCE employs a dual-path strategy, encompassing intrinsic reflection and extrinsic independent generation to correct the identified error, yielding the final, refined $CoT_{final}$.
Model Configuration and Hyperparameters All experiments were conducted on an Ubuntu 20.04 LTS server equipped with four NVIDIA GeForce RTX 3090 GPUs (24GB each), using PyTorch 1.13.0. We fine‑tuned each base LLM via Low‑Rank Adaptation (LoRA) Hu et al. (2022) with rank $r=8$ and scale factor $\alpha=16$. We utilize LLMLingua-2 Pan et al. (2024) as the token importance metric to generate our compressed CoT training data. IRM is evaluated over discrete compression rates in the range $\{0.5,0.6,0.7,0.8,0.9,1.0\}$. The $I(k)=w_{a}\cdot e^{\alpha(k-1)}$ was calibrated by fitting to data from controlled error‑injection trials. All fine‑tuning used best‑practice hyperparameters. We train the models for 5 epochs on both datasets. The peak learning rate is set to 5e55e-5, following a cosine decay schedule. We use AdamW Loshchilov and Hutter (2017) for optimization, with a warmup ratio of 0.1. We implement our training process using the LLaMA-Factory Zheng et al. (2024) library. During inference, the maximum number of tokens set to 512 for GSM8K and 1024 for MATH.
A.2. Error‑Injection Experiments
A.2.1. Error Simulation
To probe the model’s resilience in a realistic manner, our error-injection methodology does not introduce arbitrary flaws. Instead, it is designed to simulate the spontaneous reasoning failures we observed in the model’s own unguided generations, thereby ensuring the ecological validity of our experiment. We focused on two distinct classes of these naturally occurring failure patterns.
The first, Numeric Errors, simulates arithmetic miscalculations. The injection process for this error type involves identifying a numerical value in a correct reasoning step and slightly altering it, for instance, by adding or subtracting a small, fixed constant value to mimic a common calculation mistake. The second, Symbolic Errors, simulates logical inconsistencies. This is achieved by identifying a primary arithmetic operator within a step and systematically replacing it with its mathematical inverse (e.g., + is replaced by -, and * by ÷). To maintain procedural integrity, we preemptively implemented constraints to avoid invalid operations, such as division by zero.
Our overall experimental procedure involved systematically injecting one such controlled error at a specific position k within a baseline of validated, correct reasoning chains. Following this injection, the model was prompted to resume the reasoning process from that point, and the accuracy of the final answer was recorded to precisely measure the positional impact of the error.
A.2.2. Distribution of Results
Table 5 illustrates the proportion of total steps across all problems. We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
When injecting a numerical error, the accuracy drops are relatively uniform across positions: at step 2 yields a drop of 59.92%, at step 3 a drop of 60.47%, and at step 4 a drop of 61.70%a miscalculation. In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points. Replacement of an operator in the first error step (2/2) reduces the accuracy by 69.67%, while an error in the final step (4/4) inflicts a reduction of 80.80%, with symbolic faults becoming increasingly catastrophic when introduced deeper into the chain.
Infer Steps Counts Accuracy
2 786 98.85
3 2125 96.19
4 2022 95.75
5 1180 94.07
6 484 90.91
7 137 85.40
8 58 67.24
9 7 57.14
Table 5. Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
A.3. Analysis of Late‑Stage Fragility
Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer. This is facilitated by the model maintaining higher semantic entropy and contextual flexibility during the initial stages of reasoning. However, as the reasoning chain progresses, the model increasingly commits to a specific semantic trajectory, gradually narrowing its generative space and becoming less sensitive to alternative interpretations. We refer to this phenomenon as semantic commitment, a state in which the model’s outputs become rigidly anchored to previously generated content. At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in disproportionately high failure rates.
Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it. Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer. However, our analysis reveals that the impact of reasoning errors is not uniform. We also find that late-stage reasoning steps are disproportionately sensitive to localized perturbations. Errors introduced near the conclusion of reasoning are much less likely to be corrected. In contrast, early-stage errors may still trigger the model’s latent self-correction mechanisms. This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
Figure 7. Example generations by ASCoT, using Qwen2.5-Instruct 7B.
A.4. Generation Examples
To provide a qualitative illustration of our method’s end-to-end functionality, we present representative examples of reasoning chains after processing by the ASCoT pipeline. The Figure 7 displays the corrected output after being processed by ASCoT. This transformation from a flawed to a correct reasoning chain highlights the core capabilities of our system. It demonstrates the AVM module successfully identifying a high-risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair. This example visually confirm ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
A.5. Future Work
We propose several promising future directions. A key avenue is the augmentation of MSCE module through integration with external, deterministic oracles like calculators, code interpreters, or knowledge-base APIs. By querying these verifiable tools, the correction engine could access ground-truth feedback, significantly enhancing its reliability and moving beyond the constraints of the model’s internal knowledge. Another important direction involves developing online adaptive learning mechanisms. Such a system would enable ASCoT to continuously self-supervise and refine its performance in real time across various domains, evolving from a static framework into a more dynamic and robust reasoning system.